# Amplitude Expansion Phase Field Crystal (APFC) Modeling based Efficient Dislocation Simulations using Fourier Pseudospectral Method

### v2

@@ Appendix B Auxiliary Results
$\left(\frac{\delta F}{\delta\eta_{j}^{*}},\phi_{j}\right)=\left(\left[B_{0}+B_{x}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla)^{2}+3v(A^{2}-|\eta_{j}|^{2})\right]\eta_{j},\phi_{j}\right)+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right).$
@@
$\frac{\delta F}{\delta\eta_{j}^{*}}$ $=\left[B_{0}+B_{x}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla)^{2}+3v(A^{2}-|\eta_{j}|^{2})\right]\eta_{j}+\frac{\delta f^{s}}{\delta\eta_{j}^{*}}.$ (B.2)
B.1.1 The reciprocal-space vectors and ${\delta f^{s}}/{\delta\eta_{j}^{*}}$ defined in the Example 2.3
In Example 2.3, the reciprocal-space vectors used to characterize the crystal symmetry and $\frac{\delta f^{s}}{\delta\eta_{j}^{*}}$ are provided according to the selected lattice symmetry as follows:
@@
    For two-dimensional triangular lattices with $N=3$, the reciprocal-space vectors are
@@
    The complex polynomial functional in (2.14) is given by
    $f^{s}_{\rm{tri}}=-2\tau\left(\eta_{1}\eta_{2}\eta_{3}+{\rm c.c.}\right).$
    Following the derivation process of $\frac{\delta F}{\delta\eta_{j}^{*}}$, we take $j=1$ as an example:
    $\left(\frac{\delta f^{s}_{\rm tri}}{\delta\eta_{1}^{*}},\phi_{1}\right)$ $=\frac{d}{dt}f^{\rm tri}(\eta_{1},\eta_{2},\eta_{3},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\eta_{3}^{*})\bigg{|}_{t=0}$
    $=\lim_{t\to 0}\frac{f^{\rm tri}(\eta_{1},\eta_{2},\eta_{3},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\eta_{3}^{*})-f^{\rm tri}(\eta_{1},\eta_{2},\eta_{3},\eta_{1}^{*},\eta_{2}^{*},\eta_{3}^{*})}{t}$
    $=\lim_{t\to 0}-\frac{2\tau}{t}\frac{1}{|\Omega|}\int_{\Omega}\left(\eta_{1}\eta_{2}\eta_{3}+(\eta_{1}^{*}+t\phi_{1})\eta_{2}^{*}\eta_{3}^{*}-\eta_{1}\eta_{2}\eta_{3}-\eta_{1}^{*}\eta_{2}^{*}\eta_{3}^{*}\right)\,{\rm d}\bm{r}$
    $=\lim_{t\to 0}-\frac{2\tau}{t}\frac{1}{|\Omega|}\int_{\Omega}\eta_{2}^{*}\eta_{3}^{*}t\phi_{1}\,{\rm d}\bm{r}$
    $=\left(-2\tau\eta_{2}^{*}\eta_{3}^{*},\phi_{1}\right).$
@@
    $\frac{\delta f^{\rm tri}}{\delta\eta_{1}^{*}}=-2\tau\eta_{2}^{*}\eta_{3}^{*}.$
@@
    $\frac{\delta f^{s}_{\rm tri}}{\delta\eta_{j}^{*}}=-2\tau\prod_{i\neq j}^{3}\eta_{i}^{*},\qquad j=1,2,3.$ (B.3)
@@
    For the body-centered cubic (BCC) lattice with $N=6$, the reciprocal-space vectors are
    $\bm{k_{1}}=(1,1,0),\quad\bm{k_{2}}=(1,0,1),\quad\bm{k_{3}}=(0,1,1),$
    $\bm{k_{4}}=(0,1,-1),\quad\bm{k_{5}}=(1,-1,0),\quad\bm{k_{6}}=(-1,0,1).$
    The complex polynomial functional in (2.14) is given by
    $f^{s}_{\rm bcc}=-$ $2\tau\left(\eta_{1}^{*}\eta_{2}\eta_{3}+\eta_{2}^{*}\eta_{3}\eta_{5}+\eta_{3}^{*}\eta_{1}\eta_{6}+\eta_{4}^{*}\eta_{5}^{*}\eta_{6}^{*}+{\rm c.c.}\right)$
@@
    Following a similar derivation process for $\frac{\delta F}{\delta\eta_{j}^{*}}$, we take $\eta_{1}^{*}$ as an example:
    $\left(\frac{\delta f^{\rm bcc}}{\delta\eta_{1}^{*}},\phi_{1}\right)$ $=\frac{d}{dt}f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\ldots,\eta_{6}^{*})\bigg{|}_{t=0}$
    $=\lim_{t\to 0}\frac{f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\ldots,\eta_{6}^{*})-f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\eta_{2}^{*},\ldots,\eta_{6}^{*})}{t}$
    $=\lim_{t\to 0}-\frac{2\tau}{t}\frac{1}{|\Omega|}\int_{\Omega}\left(\eta_{2}\eta_{4}t\phi_{1}+\eta_{3}\eta_{6}^{*}t\phi_{1}\right)\,{\rm d}\bm{r}$
    $\quad+\lim_{t\to 0}\frac{6v}{t}\frac{1}{|\Omega|}\int_{\Omega}\left(\eta_{2}\eta_{5}^{*}\eta_{6}^{*}t\phi_{1}+\eta_{3}\eta_{4}\eta_{5}t\phi_{1}\right)\,{\rm d}\bm{r}$
    $=\frac{1}{|\Omega|}\int_{\Omega}\left(-2\tau(\eta_{2}\eta_{4}+\eta_{3}\eta_{6}^{*})+6v(\eta_{2}\eta_{5}^{*}\eta_{6}^{*}+\eta_{3}\eta_{4}\eta_{5})\right)\phi_{1}\,{\rm d}\bm{r}.$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{1}^{*}}=-2\tau(\eta_{2}\eta_{4}+\eta_{3}\eta_{6}^{*})+6v(\eta_{2}\eta_{5}^{*}\eta_{6}^{*}+\eta_{3}\eta_{4}\eta_{5}).$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{i}^{*}}=-2\tau(\eta_{k}\eta_{n}^{*}+\eta_{j}\eta_{l})+6v(\eta_{k}\eta_{l}\eta_{m}+\eta_{j}\eta_{m}^{*}\eta_{n}^{*}).$ (B.4)
@@
    $\left(\frac{\delta f^{\rm bcc}}{\delta\eta_{4}^{*}},\phi_{4}\right)$ $=\frac{d}{dt}f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\ldots,\eta_{4}^{*}+t\phi_{4},\ldots,\eta_{6}^{*})\bigg{|}_{t=0}$
    $=\lim_{t\to 0}\frac{f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\ldots,\eta_{4}^{*}+t\phi_{4},\ldots,\eta_{6}^{*})-f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\eta_{2}^{*},\ldots,\eta_{6}^{*})}{t}$
    $=\frac{1}{|\Omega|}\int_{\Omega}\left(-2\tau(\eta_{5}^{*}\eta_{6}^{*}\phi_{4}+\eta_{1}\eta_{2}^{*}\phi_{4})+6v(\eta_{1}\eta_{3}^{*}\eta_{5}^{*}\phi_{4}+\eta_{3}\eta_{2}^{*}\eta_{6}^{*}\phi_{4})\right)\,{\rm d}\bm{r}$
    $=\left(-2\tau(\eta_{5}^{*}\eta_{6}^{*}+\eta_{1}\eta_{2}^{*})+6v(\eta_{1}\eta_{3}^{*}\eta_{5}^{*}+\eta_{3}\eta_{2}^{*}\eta_{6}^{*}),\phi_{4}\right).$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{4}^{*}}=-2\tau(\eta_{5}^{*}\eta_{6}^{*}+\eta_{1}\eta_{2}^{*})+6v(\eta_{1}\eta_{3}^{*}\eta_{5}^{*}+\eta_{3}\eta_{2}^{*}\eta_{6}^{*}).$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{l}^{*}}=-2\tau(\eta_{m}^{*}\eta_{n}^{*}+\eta_{i}\eta_{j}^{*})+6v(\eta_{i}\eta_{k}^{*}\eta_{m}^{*}+\eta_{k}\eta_{j}^{*}\eta_{n}^{*}).$ (B.5)
    In this case, all the equations for the amplitudes in (B.4) and (B.5) are obtained by permutations of the groups $(i,j,k)=(1,2,3)$ and $(l,m,n)=(4,5,6)$.
@@ B.2 The calculation of $\phi_{0}$
Let $\phi_{j}=\phi_{0}$ represent the amplitude of the equilibrium crystal. By assuming the amplitudes for all $\eta_{j}$ to be real and equal, we can set $\eta_{j}=\phi_{0}$. This can be determined by minimizing the energy functional given in (2.15).
@@
    For two-dimensional triangular lattices with $N=3$,
    $f^{s}=f^{s}_{\rm{tri}}=-2\tau(\eta_{1}\eta_{2}\eta_{3}+{\rm c.c.}).$
    At this point, $F_{\eta_{j}}$ defined by (B.1) becomes
    $F_{\eta_{j}}=\frac{1}{|\Omega|}\int_{\Omega}\left(\frac{45}{2}\phi_{0}^{4}-4\tau\phi_{0}^{3}+3B_{0}\phi_{0}^{2}\right)\,{\rm d}\bm{r}.$ (B.6)
    According to the mean value theorem, the minimum value is obtained when
    $\phi_{0}=\frac{\tau+\sqrt{\tau^{2}-15\nu B_{0}}}{15v}.$ (B.7)
@@
    For the body-centered cubic (BCC) lattice with $N=6$,
    $f^{s}=f^{s}_{\rm bcc}=-$ $2\tau(\eta_{1}^{*}\eta_{2}\eta_{3}+\eta_{2}^{*}\eta_{3}\eta_{5}+\eta_{3}^{*}\eta_{1}\eta_{6}+\eta_{4}^{*}\eta_{5}^{*}\eta_{6}^{*}+{\rm c.c.})$
@@
    At this point, $F_{\eta_{j}}$ defined by (B.1) becomes
    $F_{\eta_{j}}=\frac{1}{|\Omega|}\int_{\Omega}\left(135\phi_{0}^{4}-16\tau\phi_{0}^{3}+6B_{0}\phi_{0}^{2}\right)\,{\rm d}\bm{r}.$ (B.8)
    According to the mean value theorem [52], the minimum value is obtained when
    $\phi_{0}=\frac{2\tau+\sqrt{4\tau^{2}-45vB_{0}}}{45v}.$ (B.9)

### v3

@@ B.1 The derivation of (14)
$\left(\frac{\delta F}{\delta\eta_{j}^{*}},\phi_{j}\right)$ $=\left(\left[B_{0}+B_{x}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla)^{2}+3v(A^{2}-\lvert\eta_{j}\rvert^{2})\right]\eta_{j},\phi_{j}\right)+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right).$
@@
$\frac{\delta F}{\delta\eta_{j}^{*}}$ $=\left[B_{0}+B_{x}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla)^{2}+3v(A^{2}-\lvert\eta_{j}\rvert^{2})\right]\eta_{j}+\frac{\delta f^{s}}{\delta\eta_{j}^{*}}.$ (51)
B.1.1 The reciprocal-space vectors and ${\delta f^{s}}/{\delta\eta_{j}^{*}}$ defined in the Example 1
In Example 1, the reciprocal-space vectors used to characterize the crystal symmetry and $\frac{\delta f^{s}}{\delta\eta_{j}^{*}}$ are provided according to the selected lattice symmetry as follows:
@@ B.1.1 The reciprocal-space vectors and ${\delta f^{s}}/{\delta\eta_{j}^{*}}$ defined in the Example 1
    For two-dimensional triangular lattices with $M=3$, the reciprocal-space vectors are
@@
    The complex polynomial functional in (14) is given by
    $f^{s}_{\rm{tri}}=-2\gamma\left(\eta_{1}\eta_{2}\eta_{3}+{\rm c.c.}\right).$
    Following the derivation process of $\delta F/\delta\eta_{j}^{*}$, we take $j=1$ as an example:
    $\left(\frac{\delta f^{s}_{\rm tri}}{\delta\eta_{1}^{*}},\phi_{1}\right)$ $=\frac{d}{dt}f^{\rm tri}(\eta_{1},\eta_{2},\eta_{3},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\eta_{3}^{*})\bigg\lvert_{t=0}$
    $=\lim_{t\to 0}\frac{f^{\rm tri}(\eta_{1},\eta_{2},\eta_{3},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\eta_{3}^{*})}{t}-\lim_{t\to 0}\frac{f^{\rm tri}(\eta_{1},\eta_{2},\eta_{3},\eta_{1}^{*},\eta_{2}^{*},\eta_{3}^{*})}{t}$
    $=\lim_{t\to 0}-\frac{2\gamma}{t}\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\left((\eta_{1}^{*}+t\phi_{1})\eta_{2}^{*}\eta_{3}^{*}-\eta_{1}^{*}\eta_{2}^{*}\eta_{3}^{*}\right)\,{\rm d}\bm{r}$
    $=\lim_{t\to 0}-\frac{2\gamma}{t}\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\eta_{2}^{*}\eta_{3}^{*}t\phi_{1}\,{\rm d}\bm{r}$
    $=\left(-2\gamma\eta_{2}^{*}\eta_{3}^{*},\phi_{1}\right).$
@@
    $\frac{\delta f^{\rm tri}}{\delta\eta_{1}^{*}}=-2\gamma\eta_{2}^{*}\eta_{3}^{*}.$
@@
    $\frac{\delta f^{s}_{\rm tri}}{\delta\eta_{j}^{*}}=-2\gamma\prod_{i\neq j}^{3}\eta_{i}^{*},\qquad j=1,2,3.$ (52)
@@
    For the body-centered cubic (BCC) lattice with $M=6$, the reciprocal-space vectors are
    $\bm{k}_{1}=(1,1,0),$ $\bm{k}_{2}$ $=(1,0,1),$ $\bm{k}_{3}$ $=(0,1,1),$
    $\bm{k}_{4}=(0,1,-1),$ $\bm{k}_{5}$ $=(1,-1,0),$ $\bm{k}_{6}$ $=(-1,0,1).$
    The complex polynomial functional in (14) is given by
    $f^{s}_{\rm bcc}=$- $2\gamma\left(\eta_{1}^{*}\eta_{2}\eta_{3}+\eta_{2}^{*}\eta_{3}\eta_{5}+\eta_{3}^{*}\eta_{1}\eta_{6}+\eta_{4}^{*}\eta_{5}^{*}\eta_{6}^{*}+{\rm c.c.}\right)$
@@
    Following a similar derivation process for $\delta F/\delta\eta_{j}^{*}$, we take $\eta_{1}^{*}$ as an example:
    $\left(\frac{\delta f^{\rm bcc}}{\delta\eta_{1}^{*}},\phi_{1}\right)$ $=\frac{d}{dt}f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\ldots,\eta_{6}^{*})\bigg\lvert_{t=0}$
    $=\lim_{t\to 0}\frac{f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*}+t\phi_{1},\eta_{2}^{*},\ldots,\eta_{6}^{*})}{t}$
    $\quad-\lim_{t\to 0}\frac{f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\eta_{2}^{*},\ldots,\eta_{6}^{*})}{t}$
    $=\lim_{t\to 0}-\frac{2\gamma}{t}\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\left(\eta_{2}\eta_{4}t\phi_{1}+\eta_{3}\eta_{6}^{*}t\phi_{1}\right)\,{\rm d}\bm{r}$
    $\quad+\lim_{t\to 0}\frac{6v}{t}\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\left(\eta_{2}\eta_{5}^{*}\eta_{6}^{*}t\phi_{1}+\eta_{3}\eta_{4}\eta_{5}t\phi_{1}\right)\,{\rm d}\bm{r}$
    $=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\left(-2\gamma(\eta_{2}\eta_{4}+\eta_{3}\eta_{6}^{*})+6v(\eta_{2}\eta_{5}^{*}\eta_{6}^{*}+\eta_{3}\eta_{4}\eta_{5})\right)\phi_{1}\,{\rm d}\bm{r}.$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{1}^{*}}=-2\gamma(\eta_{2}\eta_{4}+\eta_{3}\eta_{6}^{*})+6v(\eta_{2}\eta_{5}^{*}\eta_{6}^{*}+\eta_{3}\eta_{4}\eta_{5}).$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{i}^{*}}=-2\gamma(\eta_{k}\eta_{n}^{*}+\eta_{j}\eta_{l})+6v(\eta_{k}\eta_{l}\eta_{m}+\eta_{j}\eta_{m}^{*}\eta_{n}^{*}).$ (53)
@@
    $\left(\frac{\delta f^{\rm bcc}}{\delta\eta_{4}^{*}},\phi_{4}\right)$ $=\frac{d}{dt}f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\ldots,\eta_{4}^{*}+t\phi_{4},\ldots,\eta_{6}^{*})\bigg\lvert_{t=0}$
    $=\lim_{t\to 0}\frac{f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\ldots,\eta_{4}^{*}+t\phi_{4},\ldots,\eta_{6}^{*})}{t}$
    $\quad-\lim_{t\to 0}\frac{f^{\rm bcc}(\eta_{1},\eta_{2},\ldots,\eta_{6},\eta_{1}^{*},\eta_{2}^{*},\ldots,\eta_{6}^{*})}{t}$
    $=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}-2\gamma(\eta_{5}^{*}\eta_{6}^{*}\phi_{4}+\eta_{1}\eta_{2}^{*}\phi_{4})\,{\rm d}\bm{r}$
    $\quad+\frac{1}{\lvert\Omega\rvert}\int_{\Omega}6v(\eta_{1}\eta_{3}^{*}\eta_{5}^{*}\phi_{4}+\eta_{3}\eta_{2}^{*}\eta_{6}^{*}\phi_{4})\,{\rm d}\bm{r}$
    $=\left(-2\gamma(\eta_{5}^{*}\eta_{6}^{*}+\eta_{1}\eta_{2}^{*})+6v(\eta_{1}\eta_{3}^{*}\eta_{5}^{*}+\eta_{3}\eta_{2}^{*}\eta_{6}^{*}),\phi_{4}\right).$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{4}^{*}}=-2\gamma(\eta_{5}^{*}\eta_{6}^{*}+\eta_{1}\eta_{2}^{*})+6v(\eta_{1}\eta_{3}^{*}\eta_{5}^{*}+\eta_{3}\eta_{2}^{*}\eta_{6}^{*}).$
@@
    $\frac{\delta f^{\rm bcc}}{\delta\eta_{l}^{*}}=-2\gamma(\eta_{m}^{*}\eta_{n}^{*}+\eta_{i}\eta_{j}^{*})+6v(\eta_{i}\eta_{k}^{*}\eta_{m}^{*}+\eta_{k}\eta_{j}^{*}\eta_{n}^{*}).$ (54)
    In this case, all the equations for the amplitudes in (53) and (54) are obtained by permutations of the groups $(i,j,k)=(1,2,3)$ and $(l,m,n)=(4,5,6)$.
@@ B.2 The calculation of $\phi_{0}$
Let $\phi_{j}=\phi_{0}$ represent the amplitude of the equilibrium crystal. By assuming the amplitudes for all $\eta_{j}$ to be real and equal, we can set $\eta_{j}=\phi_{0}$. This can be determined by minimizing the energy functional given in (15).
@@
    For two-dimensional triangular lattices with $M=3$,
    $f^{s}=f^{s}_{\rm{tri}}=-2\gamma(\eta_{1}\eta_{2}\eta_{3}+{\rm c.c.}).$
    At this point, $F$ defined by (50) becomes
    $F=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\left(\frac{45}{2}v\phi_{0}^{4}-4\gamma\phi_{0}^{3}+3B_{0}\phi_{0}^{2}\right)\,{\rm d}\bm{r}.$ (55)
    According to the necessary condition for extrema of differentiable functions, the minimum value is obtained when
    $\phi_{0}=\frac{\gamma+\sqrt{\gamma^{2}-15vB_{0}}}{15v}.$ (56)
@@
    For the body-centered cubic (BCC) lattice with $M=6$,
    $f^{s}=f^{s}_{\rm bcc}=$- $2\gamma(\eta_{1}^{*}\eta_{2}\eta_{3}+\eta_{2}^{*}\eta_{3}\eta_{5}+\eta_{3}^{*}\eta_{1}\eta_{6}+\eta_{4}^{*}\eta_{5}^{*}\eta_{6}^{*}+{\rm c.c.})$
@@
    At this point, $F$ defined by (50) becomes
    $F=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\left(135v\phi_{0}^{4}-16\gamma\phi_{0}^{3}+6B_{0}\phi_{0}^{2}\right)\,{\rm d}\bm{r}.$ (57)
    According to the necessary condition for extrema of differentiable functions, the minimum value is obtained when
    $\phi_{0}=\frac{2\gamma+\sqrt{4\gamma^{2}-45vB_{0}}}{45v}.$ (58)